– General Linux 1 –

Configure Fundamental BIOS Settings [3]

(Linux Professional Institute Certification)

a

   .~.
   /V\      by: geoffrey robertson
  // \\         geoffrey@zip.com.au
  @._.@
$Id: gl1.101.1.slides.tex,v 1.4 2003/05/30 04:56:33 waratah Exp $

aCopyright \begin{matrix}c \\ ◯ \end{matrix} 2002 Geoffrey Robertson. Permission is granted to make and distribute verbatim copies or modified versions of this document provided that this copyright notice and this permission notice are preserved on all copies under the terms of the GNU General Public License as published by the Free Software Foundation—either version 2 of the License or (at your option) any later version.

Configure BIOS Fundamental Settings [3]

(1.1) 101 Hardware & Architecture [11]

1.101.1

Configure BIOS Fundamental Settings [3] 1.101.3 Configure Modem and Sound Cards [2] 1.101.4 Setup SCSI Devices [2] 1.101.5 Set Up Different PC Expansion Cards [2] 1.101.6 Configure Communication Devices [2] 1.101.7 Set Up USB devices [2]

Configure BIOS Fundamental Settings [3]

Objective

Candidates should be able to configure fundamental system hardware by making the correct settings in the system BIOS. This objective includes a proper understanding of BIOS configuration issues such as the use of LBA on IDE hard disks larger than 1024 cylinders, enabling or disabling integrated peripherals, as well as configuring systems with (or without) external peripherals such as keyboards. It also includes the correct setting for IRQ, DMA and I/O addresses for all BIOS administrated ports and settings for error handling.

Configure Fundamental BIOS Settings [3]

Key files, terms, and utilities

/proc/ioports
/proc/interrupts
/proc/dma
/proc/pci

Configure Fundamental BIOS Settings [3]

Resources of interest

Large Disk HOWTO by Andries Brouwer

http://www.linuxdoc.org/HOWTO/Large-Disk-HOWTO.html

The PC Firmware—BIOS

Responsible for bring the hardware to a state at which it is ready to boot an Operating System.

• System reset and initialisation

• Basic main memory test (RAM)

• Test the presence of various devices

• On board peripheral selection

• Low level device configuration

• Selection of boot devices

• Locating an operating system

• The BIOS may have password access enabled

The PC Firmware—BIOS

Responsible for bring the hardware to a state at which it is ready to boot an Operating System.

• System reset and initialisation

• Basic main memory test (RAM)

• Test the presence of various devices

• On board peripheral selection

• Low level device configuration

• Selection of boot devices

• Locating an operating system

• The BIOS may have password access enabled

6-a

The PC Firmware—BIOS

Responsible for bring the hardware to a state at which it is ready to boot an Operating System.

• System reset and initialisation

• Basic main memory test (RAM)

• Test the presence of various devices

• On board peripheral selection

• Low level device configuration

• Selection of boot devices

• Locating an operating system

• The BIOS may have password access enabled

6-b

The PC Firmware—BIOS

Responsible for bring the hardware to a state at which it is ready to boot an Operating System.

• System reset and initialisation

• Basic main memory test (RAM)

• Test the presence of various devices

• On board peripheral selection

• Low level device configuration

• Selection of boot devices

• Locating an operating system

• The BIOS may have password access enabled

6-c

The PC Firmware—BIOS

Responsible for bring the hardware to a state at which it is ready to boot an Operating System.

• System reset and initialisation

• Basic main memory test (RAM)

• Test the presence of various devices

• On board peripheral selection

• Low level device configuration

• Selection of boot devices

• Locating an operating system

• The BIOS may have password access enabled

6-d

The PC Firmware—BIOS

Responsible for bring the hardware to a state at which it is ready to boot an Operating System.

• System reset and initialisation

• Basic main memory test (RAM)

• Test the presence of various devices

• On board peripheral selection

• Low level device configuration

• Selection of boot devices

• Locating an operating system

• The BIOS may have password access enabled

6-e

The PC Firmware—BIOS

Responsible for bring the hardware to a state at which it is ready to boot an Operating System.

• System reset and initialisation

• Basic main memory test (RAM)

• Test the presence of various devices

• On board peripheral selection

• Low level device configuration

• Selection of boot devices

• Locating an operating system

• The BIOS may have password access enabled

6-f

The PC Firmware—BIOS

Responsible for bring the hardware to a state at which it is ready to boot an Operating System.

• System reset and initialisation

• Basic main memory test (RAM)

• Test the presence of various devices

• On board peripheral selection

• Low level device configuration

• Selection of boot devices

• Locating an operating system

• The BIOS may have password access enabled

6-g

The PC Firmware—BIOS

Responsible for bring the hardware to a state at which it is ready to boot an Operating System.

• System reset and initialisation

• Basic main memory test (RAM)

• Test the presence of various devices

• On board peripheral selection

• Low level device configuration

• Selection of boot devices

• Locating an operating system

• The BIOS may have password access enabled

6-h

The PC Firmware—BIOS

Accessing BIOS Setup

The various vendors have different ways of entering the BIOS menu system • Del

• F2

• Fn Setup

• Special Boot Floppy •

The PC Firmware—BIOS

Boot devices

• Floppy

• CDROM / DVD • IDE Disk

• SCSI Disk

• Network

• Zip / Jazz

• USB Disk

The PC Firmware—BIOS

Resource Assignment

• PC’s have memory mapped I/O and devices must not have overlapping I/O Addresses

• Devices sharing interrupts such as ttyS0 and ttyS2 must not be used at the same time

• Direct Memory Access (DMA) allows some devices to access memory directly, rather than via the CPU

• Some PCI controllers may have to be in a Bus Mastering Slot

• PCI devices may share a common interrupt

The PC Firmware—BIOS

Resource Assignment

• PC’s have memory mapped I/O and devices must not have overlapping I/O Addresses

• Devices sharing interrupts such as ttyS0 and ttyS2 must not be used at the same time

• Direct Memory Access (DMA) allows some devices to access memory directly, rather than via the CPU

• Some PCI controllers may have to be in a Bus Mastering Slot

• PCI devices may share a common interrupt

9-a

The PC Firmware—BIOS

Resource Assignment

• PC’s have memory mapped I/O and devices must not have overlapping I/O Addresses

• Devices sharing interrupts such as ttyS0 and ttyS2 must not be used at the same time

• Direct Memory Access (DMA) allows some devices to access memory directly, rather than via the CPU

• Some PCI controllers may have to be in a Bus Mastering Slot

• PCI devices may share a common interrupt

9-b

The PC Firmware—BIOS

Resource Assignment

• PC’s have memory mapped I/O and devices must not have overlapping I/O Addresses

• Devices sharing interrupts such as ttyS0 and ttyS2 must not be used at the same time

• Direct Memory Access (DMA) allows some devices to access memory directly, rather than via the CPU

• Some PCI controllers may have to be in a Bus Mastering Slot

• PCI devices may share a common interrupt

9-c

The PC Firmware—BIOS

Resource Assignment

• PC’s have memory mapped I/O and devices must not have overlapping I/O Addresses

• Devices sharing interrupts such as ttyS0 and ttyS2 must not be used at the same time

• Direct Memory Access (DMA) allows some devices to access memory directly, rather than via the CPU

• Some PCI controllers may have to be in a Bus Mastering Slot

• PCI devices may share a common interrupt

9-d

The PC Firmware—BIOS

Common Device Settings

Device I/O Address IRQ DMA
ttyS0  3F8  4  NA
ttyS1  2F8  3  NA
ttyS2  3E8  4  NA
ttyS3  2E8  3  NA
lp0  378-37F  7  NA
lp1  278-27F  5  NA
fd0, fd1 3F0-3F7  6  2
fd2, fd3 370-377  10  3

The PC Firmware—BIOS

1023 Cylinder Limit

• Ten bit addressing of cylinders—11 1111 1111_{2} = 1023_{10}

• Large Block Addressing (LBA) used for disks larger than 1023 cylinders (≈8Gb)

• The boot loader must be below cylinder 1024

• Older versions of LILO must have /boot (the kernel) entirely below cylinder 1024

• The Grub bootloader does not have these restrictions

The PC Firmware—BIOS

1023 Cylinder Limit

• Ten bit addressing of cylinders—11 1111 1111_{2} = 1023_{10}

• Large Block Addressing (LBA) used for disks larger than 1023 cylinders (≈8Gb)

• The boot loader must be below cylinder 1024

• Older versions of LILO must have /boot (the kernel) entirely below cylinder 1024

• The Grub bootloader does not have these restrictions

11-a

The PC Firmware—BIOS

1023 Cylinder Limit

• Ten bit addressing of cylinders—11 1111 1111_{2} = 1023_{10}

• Large Block Addressing (LBA) used for disks larger than 1023 cylinders (≈8Gb)

• The boot loader must be below cylinder 1024

• Older versions of LILO must have /boot (the kernel) entirely below cylinder 1024

• The Grub bootloader does not have these restrictions

11-b

The PC Firmware—BIOS

1023 Cylinder Limit

• Ten bit addressing of cylinders—11 1111 1111_{2} = 1023_{10}

• Large Block Addressing (LBA) used for disks larger than 1023 cylinders (≈8Gb)

• The boot loader must be below cylinder 1024

• Older versions of LILO must have /boot (the kernel) entirely below cylinder 1024

• The Grub bootloader does not have these restrictions

11-c

The PC Firmware—BIOS

1023 Cylinder Limit

• Ten bit addressing of cylinders—11 1111 1111_{2} = 1023_{10}

• Large Block Addressing (LBA) used for disks larger than 1023 cylinders (≈8Gb)

• The boot loader must be below cylinder 1024

• Older versions of LILO must have /boot (the kernel) entirely below cylinder 1024

• The Grub bootloader does not have these restrictions

11-d

            Viewing /proc/*
 /proc/dma
$ cat /proc/dma ↩
4: cascade

 /proc/interrupts
$ cat /proc/interrupts ↩
             CPU0
  0:     2623139           XT-PIC  timer
  1:       11619           XT-PIC  keyboard
  2:           0           XT-PIC  cascade
  5:       73992           XT-PIC  Maestro3(i)
  8:           1           XT-PIC  rtc
 11:         308           XT-PIC  usb-uhci,
Texas Instruments PCI4451 PC card Cardbus Controller,
Texas Instruments PCI4451 PC card Cardbus Controller, eth0
 12:      926438           XT-PIC  PS/2 Mouse
 14:      504452           XT-PIC  ide0
NMI:           0
ERR:           0

 /proc/ioports
$ cat /proc/ioports ↩
 0000-001f : dma1
0020-003f : pic1
0040-005f : timer
0060-006f : keyboard
0070-007f : rtc
0080-008f : dma page reg
00a0-00bf : pic2
00c0-00df : dma2
00f0-00ff : fpu
01f0-01f7 : ide0
02f8-02ff : serial(auto)
0378-037a : parport0
03c0-03df : vga+

/proc/pci
$ cat /proc/pci ↩
PCI devices found:
Bus  0, device   0, function  0:
Host bridge: Intel Corp. 82815 815 Chipset Host Bridge
                        and Memory Controller Hub (rev 2).
Prefetchable 32 bit memory at 0xe4000000 [0xe7ffffff].
Bus  0, device   1, function  0:
PCI bridge: PCI device 8086:1131 (Intel Corp.) (rev 2).
  Master Capable.  Latency=32.  Min Gnt=12.
Bus  0, device  30, function  0:
PCI bridge: Intel Corp. 82820 820 (Camino 2) Chipset
                                        PCI (-M) (rev 3).
  Master Capable.  No bursts.  Min Gnt=6.

The End

16